# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

### v1

@@ III-A1 Min strategy
where ||$\cdot$|| ⋅ | represents the cardinality of a set.
The filtering procedure consists of two steps: (1) selecting a suitable frequency cutoff, and (2) adjusting the grouping results to satisfy the min strategy defined in Definition 2. According to Definition 2, for fixed local centers, the grouping result satisfying the max strategy is unique, whereas multiple solutions may satisfy the min strategy. Thus, a reasonable initial grouping method that provides high-quality results and estimates the number of groups is essential. We summarize the designed grouping method in Algorithm 1. This method leverages local depths and the similarities between neighborhoods of local centers. Here, local depth values are computed during defining local centers, and the similarity matrix for neighborhoods of local centers is denoted as $\mathbb{M}$, with $\mathrm{sim}(\mathbf{N}_{c_{i}},\mathbf{N}_{c_{j}})$ as its $ij$-th entry.
Algorithm 1 Local depth and similarity-based grouping
1:Local centers $c_{1},\ldots,c_{T}$, corresponding top $s/2$ neighbors, local depth values (LD), similarity matrix $\mathbb{M}$, and a similarity threshold range $[0.4,0.75]$.
2:Grouping results $\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{G}\}$.
3:For each local center $c_{i}$, set valid neighbors $\mathrm{VN}_{i}$ as local centers in its top $s/2$ neighbors (excluding itself).
4:Initialize mean similarities and closest neighbors: MS=zero array of size $\mathrm{MS}=\text{zero array of size }T$, $\mathrm{CN}=1{:}T$
5:for $t=1$ to $T$ do
6:     Compute mean similarity: $\mathrm{MS}_{t}=\frac{\sum_{c_{j}\in\mathrm{VN}_{t}}\mathbb{M}_{tj}}{|\mathrm{VN}_{t}|}$.
7:     Find the closest neighbor: $\mathrm{CN}_{t}=\arg\max_{j\neq t}\mathbb{M}_{tj}$.
8:end for
9:Set similarity threshold $\delta_{i}$ for each local center $c_{i}$:
$\delta_{i}=\min(\mathrm{MS}_{i},\mathrm{MS}_{\mathrm{CN}_{i}})$
and if $\delta_{i}\notin[0.4,0.75]$ adjust it to the nearest bound.
10:Initialize stable centers $\mathrm{SC}=\emptyset$.
11:for $t=1$ to $T$ do
12:     if All $c_{j}\in\mathrm{VN}_{t}$ with $\mathrm{sim}(\mathbf{N}_{c_{t}},\mathbf{N}_{c})\geq\delta_{t}$ have $\text{LD}_{j}<\text{LD}_{t}$ then
13:         Add $c_{t}$ to $\mathrm{SC}$.
@@
15:end for
16:Set $G=|\mathrm{SC}|$.
17:Assign each remaining local center to the group of its most similar stable center using $\mathbb{M}$.
18:Return groups $\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{G}\}$.
Algorithm 1 uses local depth to assess the centrality of each center and employs neighborhood similarity to evaluate the relationship between centers. The similarity threshold is determined automatically, eliminating the need for manual parameter selection. Specifically, the threshold for each center $c$ is set as the minimum of the mean similarity between $c$ and its closest neighbor. This approach provides a center-specific threshold while ensuring that centers with similar neighborhoods share comparable thresholds. The empirical threshold range $[0.4,0.75]$ prevents the values from being excessively small or large. Moreover, the algorithm naturally estimates the number of groups, which corresponds to the number of estimated clusters.
A limitation of Algorithm 1 is if multiple clusters overlap with each other, the local center defined in the overlapped region may have high local depth values. These local centers can be removed by analyzing the number of unique neighbors (Definition 3) within each group, as groups containing local centers from overlapped regions tend to have very few unique neighbors. Specifically, we consider a group of centers should be dropped if the number of unique neighbors of this group is less than $s/10$. Additionally, local centers highly similar to those in the identified group should also be dropped to avoid ambiguity.
With the grouping of local centers, we determine candidates for the cutoff in the order of frequency. The first candidate index is defined as
$\max\left(\arg\min_{t}\{\mathcal{P}_{t}\geq 0.75\},\max_{\mathbf{g}}\{\arg\min_{t:c_{t}\in\mathbf{g}}t\}\right),$ (12)
where the first term ensures the final cumulative proportion $\mathcal{P}_{t}$ exceeds $0.75$ to prevent excessive elimination of local centers, and the second term guarantees at least one local center from each group is retained.
Additional candidates include local centers after the begin index that provide complementary information beyond those in the same group with higher frequency. Specifically, a candidate index $i$ must satisfy two conditions: First for all $c_{j}$ where $j\leq i$ and $c_{j}$ belongs to the same group $\mathbf{g}$ as $c_{i}$,
$\sum_{j}\mathbb{M}_{ij}=\min_{l\leq i,c_{l}\in\mathbf{g}}\sum_{j}\mathbb{M}_{lj},$ (13)
This ensures $c_{i}$ is the least similar to other local centers with higher frequencies in its group. Second, for all $c_{v}$ where $v<i$ and $c_{v}\notin\mathbf{g}$,
$\max_{v}\mathbb{M}_{iv}\leq\min_{j<i,c_{j}\in\mathbf{g}}\max_{v}\mathbb{M}_{jv},$ (14)
ensuring $c_{i}$ is less similar to points in other groups compared to any existing local center in its own group. In essence, these candidates are distant from local centers in other groups while also comparatively distinct within their own group, making them potential sources of complementary information.
Lastly, we evaluate each candidate in ascending order and count the number of “bad points” between candidates, defined as local centers that fail to satisfy Equations (9) or (10) after removing local centers in groups with indices larger than the candidate. If the number of bad points between two consecutive candidates $c_{i}$ and $c_{i^{\prime}}$ exceeds $1$, we stop at $c_{i}$.
In the above setting, the initial candidate should have a cumulative proportion of at least $0.75$. However, if $\mathcal{P}_{T}<0.75$, it may indicate that overlapping clusters cause local centers to concentrate in shared regions rather than capturing true subcluster centers. The robustness of the depth median (e.g. SD’s breakdown point is $0.5$) biases local centers toward overlapping regions when most points in a cluster have neighborhoods with similarities exceeding $0.5$ to points in other clusters. To address this, if $\mathcal{P}_{T}<0.75$, we retain only centers dissimilar to those in other groups after grouping $c_{1},\ldots,c_{T}$ using Algorithm 1. For $c\in\mathbf{g}$, we check:
$\frac{\sum_{j:c_{j}\in\mathbf{g}\setminus c_{i}}\mathbb{M}_{ij}}{|\mathbf{g}|-1}-\max_{v:c_{v}\notin\mathbf{g}}\mathbb{M}_{iv}>0.$ (15)
Then, we iteratively add back local centers from the dropped ones, including those with $f=1$ or $r_{i}(\mathbf{x}_{i})>2$. In each step, the local center that maximizes the increase in cumulative proportion is added back. The process stops when the cumulative proportion exceeds $90\%$ of the cumulative proportion of all local centers.
Algorithm 2 Local center filtering under the min strategy
1:All local centers, their ranks, frequencies, local depth values, neighborhoods, size $s$ and similarity matrix $\mathbb{M}$.
2:Filtered local centers and their local depth values.
3:Perform initial filtering to obtain $\{c_{1},\ldots,c_{T}\}$.
4:if $\mathcal{P}_{T}\geq 0.75$ then
5:     Apply Algorithm 1 to obtain $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$.
6:     Set Continue as True.
7:     while Continue do
8:         Obtain unique neighbors $\{\mathbf{U}_{\mathbf{g}_{1}},\ldots,\mathbf{U}_{\mathbf{g}_{G}}\}$.
9:         Identify $\mathrm{di}=\arg\min_{i:|\mathbf{U}_{\mathbf{g}_{i}}|<s/10}|\mathbf{U}_{\mathbf{g}_{i}}|$.
10:         if $\mathrm{di}\neq\emptyset$ then
11:              Set $\delta=\max_{i,j\in\mathbf{g}_{\mathrm{di}},i\neq j}\mathbb{M}_{ij}$.
12:              Mark centers in $\mathbf{g}_{\mathrm{di}}$ as drop centers.
13:              Add centers $v$ outside $\mathbf{g}_{\mathrm{di}}$ to drop centers if $\max_{i:c_{i}\in\mathbf{g}_{\mathrm{di}}}\mathbb{M}_{vi}\geq\delta$.
14:              Remove $\mathbf{g}_{\mathrm{di}}$ and drop centers, and update $G$.
15:         else
16:              Set Continue as False;
17:         end if
18:     end while
19:     Define the first cutoff candidate $\mathrm{cc}_{1}$ as in (12).
20:     Identify additional candidates among centers with indices greater than $\mathrm{cc}_{1}$ and satisfy (13) and (14).
21:     Initialize bp=zero array of size |$\mathrm{bp}=\text{zero array of size }|\mathrm{cc}|$.
22:     for $q=1$ to $|\mathrm{cc}|$ do
23:         Remove local centers in groups with indices larger than $\mathrm{cc}_{q}$ to form $\{\mathbf{g}^{\prime}_{1},\ldots,\mathbf{g}^{\prime}_{G}\}$.
[…]
25:     end for
26:     Define cutoff as $\mathrm{cc}_{p}$ where $p$ is the first index satisfying $\mathrm{bp}_{p+1}-\mathrm{bp}_{p}>1$. If no such $p$ exists, set $cc_{p}$ as the largest cutoff candidate index.
27:else
28:     Drop local centers from $\{c_{1},\ldots,c_{T}\}$ that do not satisfy (15).
29:     Compute current cumulative proportion $\mathcal{P}$ and total cumulative proportion $\mathcal{P}^{\prime}$.
30:     while $\mathcal{P}<0.9\mathcal{P}^{\prime}$ do
31:         Add the local center maximizes the increase in $\mathcal{P}$.
32:         Update $\mathcal{P}$.
33:     end while
34:end if
35:Save remaining local centers and their local depth values.
After Algorithm 2, we reapply Algorithm 1 to group the filtered centers. If any local centers fail to meet the requirements of the min strategy, Algorithm 3 is applied, which is referred to as the group trimming process. The number of clusters is not required, but if users specify $K$, the algorithm forces the number of groups to equal $K$. For simplicity, the left and right sides of inequality (9) are denoted as $\mathrm{mw}(c)$ and $\mathrm{mb}(c)$ in the algorithm, respectively.
@@
1:Groups of local centers $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$, similarity matrix $\mathbb{M}$, optional $K$.
2:Trimmed groups satisfying the min strategy.
3:if $K\neq\emptyset$ then
4:     if $G>K$ then
5:         Drop groups with the highest $\{\arg\min_{t:c_{t}\in\mathbf{g}}t\}$ until $G$ equals to $K$.
6:     else
7:         while $G<K$ do
8:              Define a new stable center and increment $G$ by $1$triangleright Details provided in supplementary material
9:         end while
10:         Update $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$.
11:     end if
12:end if
13:while violations of (9) or (10) exist do
14:     while violations of (9) exist do
15:         for each group ${\mathbf{g}}_{l}$ with violations of (9do
16:              Drop the center $c\in{\mathbf{g}}_{l}$ which maximizes:
$\frac{\sum_{c^{\prime}\in{\mathbf{g}}_{l}\setminus\{c\}}\left(\mathrm{mw}(c^{\prime})-\mathrm{mb}(c^{\prime})\right)}{|{\mathbf{g}}_{l}\setminus\{c\}|}.$
17:         end for
18:         Update groups and recompute violations of (9).
19:     end while
20:     if violations of (10) exist then
21:         Drop all centers violating (10).
22:         Update groups and recompute violations.
23:     end if
24:end while

### v2

@@ III-A1 Min strategy
where $|\cdot|$ represents the cardinality of a set.
The filtering procedure under the min strategy consists of two stages: (1) selecting a frequency cutoff and forming an grouping, and (2) refining the grouping to satisfy the min strategy in Definition 2. For a fixed set of local centers, the max strategy yields a unique grouping, whereas the min strategy may admit multiple feasible groupings. Thus, we use a heuristic procedure that performs filtering and grouping simultaneously. Stage (1) is implemented by Algorithms 2 with grouping detailed in Algorithm 1, and stage (2) by the group trimming process in Algorithm 3. Implementation details are given next and may be skipped on a first reading without loss of the main ideas of the paper.
Our method leverages a stability criterion for evaluating a local center $c_{i}$, termed average neighbor-induced depth (AND). Recall $\mathbb{S}_{j,i}=D_{\mathrm{SD}}(\mathbf{x}_{i}\mid\mathbf{X}_{Rj})$ denote the depth-based similarity from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$, we define AND as
$\mathrm{AND}(\mathbf{x}_{i})\;=\;\frac{1}{s}\sum_{j\in\mathbf{N}_{i}}\mathbb{S}_{j,i}.$ (10)
Intuitively, $\mathrm{AND}(i)$ aggregates the depth of $\mathbf{x}_{i}$ computed w.r.t. reflected datasets centered at its neighbors. It tends to be large only when $\mathbf{x}_{i}$ is consistently central relative to the collection of reflected data sets induced by its neighborhood.
One might ask why local depth is not adopted here. As noted in the literature [40, 43], local depth may assign relatively high values to points located between two distributions, which can be central in a geometric sense but not necessarily lie in regions of high data concentration. In clustering tasks, such behavior may lead to the selection of misleading representative centers.
We next identify stable centers from the frequency-sorted local-center candidates. For each candidate $c_{t}$, we consider its top $\lfloor s/2\rfloor$ depth-based neighbors (i.e., half points in $\mathbf{N}_{c}$), and two candidates are regarded as mutual neighbors if they appear in each other’s top $\lfloor s/2\rfloor$ neighbors. The set of mutual neighbors of local center $c$ is denoted as $\mathrm{MN}(c)$
A candidate $c_{t}$ is declared a stable center if its AND is not smaller than that of any of its mutual neighbors. Let $\mathrm{SC}$ be the set of stable centers. We take the cutoff index as the last stable center in the frequency order and retain the local centers up to this index. To avoid premature truncation, we further check the cumulative proportion $\mathcal{P}_{t}$. If $\mathcal{P}_{t}<0.75$, we either (1) simply extend the frequency cutoff until the coverage first exceeds $0.75$ when the gap is large, or (2) when the coverage is close to $0.75$, greedily add a small number of additional candidates that maximize the product of marginal coverage gain and AND, thereby efficiently increasing coverage without introducing many low-quality centers.
Let $\mathbf{C}$ denote the retained local-center set, and let $\mathbb{M}$ collect the pairwise neighborhood similarities among centers in $\mathbf{C}$, with entries $\mathbb{M}_{i,j}=\mathrm{sim}(\mathbf{N}_{c_{i}},\mathbf{N}_{c_{j}})$ computed by (6). For each stable center $c_{t}$, we form a group by aggregating centers whose similarities to $c_{t}$ exceed an adaptive threshold $\delta_{t}$. Throughout the algorithm, all such similarity thresholds are constrained to lie in $\delta_{\min},\delta_{\max}]$ and in all experiments, we set $(\delta_{\min},\delta_{\max})=(0.4,0.7)$ to avoid overly loose or overly stringent grouping.
$\delta_{t}=\max\!\Bigl\{$ $\min_{c_{u}\in\mathrm{MN}(c_{t})\cap\mathbf{C}}\mathbb{M}_{t,u}\;\text{s.t.}\;\mathbb{M}_{t,u}>\max_{c_{v}\in\mathbf{C}\setminus\mathrm{MN}(c_{t})}\mathbb{M}_{t,v},$ (11)
$\delta_{\min}\Bigr\}.$
This definition ensures that $\delta_{t}$ retains only mutual neighbors of $c_{t}$ while excluding non-mutual candidates, subject to a conservative lower bound that prevents overly loose thresholds.
The initial grouping may not cover all local centers in $\mathbf{C}$. We therefore perform a second pass on the unassigned candidates, applying the same rules for defining extra stable centers but restricting mutual-neighbor relations to the unassigned set. When a candidate center $c_{t}$ has no remaining mutual neighbors, we assess its stability by whether its neighborhood $\mathbf{N}_{c_{t}}$ contributes sufficient new coverage beyond the neighborhoods of all other local centers in $\mathbf{C}$. Specifically, we define $c_{t}$ as a stable center only if
$\left|\mathbf{N}_{c_{t}}\setminus\bigcup_{c_{u}\in\mathbf{C}\setminus\{c_{t}\}}\mathbf{N}_{c_{u}}\right|>\frac{s}{3}.$ (12)
We then use the same rule for creating groups of newly defined stable centers.
Preliminary groups formed around stable centers may overlap. We resolve such overlaps and decide possible merges using a lightweight two-stage heuristic. Consider two groups $\mathbf{g}_{p}$ and $\mathbf{g}_{q}$ with nonempty intersection (the extension to overlaps involving more than two groups is similar). We define their cores by removing the overlapped centers:
$\mathrm{core}(\mathbf{g}_{p})=\mathbf{g}_{p}\setminus(\mathbf{g}_{p}\cap\mathbf{g}_{q}),\qquad\mathrm{core}(\mathbf{g}_{q})=\mathbf{g}_{q}\setminus(\mathbf{g}_{p}\cap\mathbf{g}_{q}).$
Each overlapped center $c\in\mathbf{g}_{p}\cap\mathbf{g}_{q}$ is hard-assigned to the group with the larger minimum neighborhood similarity between $c$ and the corresponding core; if both are below $\delta_{\min}$, $c$ is discarded for ambiguity.
After this, we test whether two previous overlapped groups should be merged. We merge $\mathbf{g}_{p}$ and $\mathbf{g}_{q}$ if
$\max_{c_{u}\in\mathbf{g}_{p},\,c_{v}\in\mathbf{g}_{q}}\mathbb{M}_{u,v}\;\geq\;\min\!\{\min_{c_{u},c_{v}\in\mathbf{g}_{p}}\mathbb{M}_{c_{u},c_{v}},\;\min_{c_{u},c_{v}\in\mathbf{g}_{q}}\mathbb{M}_{c_{u},c_{v}}\},$ (13)
and
$\min_{c_{u},c_{v}\in\mathbf{g}_{p}\cup\mathbf{g}_{q}}\mathbb{M}_{u,v}\geq\delta_{\min}.$ (14)
In implementation, the within-group minima in (13) are clipped to $\delta_{\min},\delta_{\max}]$.
Lastly, if some local centers in $\mathbf{C}$ remain unassigned, we assign each such center $c_{t}$ to
$g^{*}_{t}=\arg\max_{g\in\mathcal{G}(c_{t})}\;\min_{c_{u}\in\mathbf{g}_{g}}\mathbb{M}_{t,u},$ (15)
where
$\mathcal{G}(c_{t})=\left\{g:\min_{c_{u}\in\mathbf{g}_{g}}\mathbb{M}_{t,u}\geq\min\!\left(\min_{c_{u},c_{v}\in\mathbf{g}_{g}}\mathbb{M}_{u,v},\,\delta_{\max}\right)\right\},$
otherwise $c$ is dropped from $\mathbf{C}$.
Algorithm 1 AND-based stable-center selection and grouping
1:Candidate local centers, neighborhoods $\{\mathbf{N}_{c}\}$ (size $s$), pointwise similarities $\mathbb{S}$, AND $\mathrm{AND}(\cdot)$ in (10), threshold bounds $\delta_{\min},\delta_{\max}]$, and sample size $n$.
2:Retained centers $\mathbf{C}$ and groups $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{\hat{K}}\}$.
3:Mutual-neighbor graph: for each candidate $c$, keep its top $\lfloor s/2\rfloor$ candidate neighbors under $\mathbb{S}$; define $\mathrm{MN}(c)$ as mutual neighbors.
4:Stable centers: $\mathrm{SC}\leftarrow\{c_{t}:\mathrm{AND}(c_{t})\geq\mathrm{AND}(c_{u}),\ \forall c_{u}\in\mathrm{MN}(c_{t})\}$.
5:Frequency cutoff: $t_{0}\leftarrow\max\{t:c_{t}\in\mathrm{SC}\}$; initialize $\mathbf{C}\leftarrow\{c_{1},\ldots,c_{t_{0}}\}$.
6:Compute $\mathcal{P}(\mathbf{C})=\frac{1}{n}\big|\cup_{c\in\mathbf{C}}\mathbf{N}_{c}\big|$ and if $\mathcal{P}(\mathbf{C})<0.75$ update $\mathbf{C}$ using the two-case rule in the text until $\mathcal{P}(\mathbf{C})\geq 0.75$.
7:Build neighborhood similarity matrix $\mathbb{M}$ over $\mathbf{C}$.
[…]
9:for each $c_{t}\in\mathrm{SC}$ do
10:  Compute $\delta_{t}$ by (11) and form $\mathbf{g}_{t}\leftarrow\{c_{u}\in\mathbf{C}:\mathbb{M}_{t,u}\geq\delta_{t}\}$.
11:  $\mathcal{G}\leftarrow\mathcal{G}\cup\{\mathbf{g}_{t}\}$.
12:end for
13:Second pass: $\mathcal{L}\leftarrow\mathbf{C}\setminus\bigcup_{\mathbf{g}\in\mathbf{G}}\mathbf{g}$.
14:if $\mathcal{L}\neq\emptyset$ then
15:  Obtain $\mathrm{SC}_{2}$ by repeating Line 2 on $\mathcal{L}$ (mutual neighbors are defined within $\mathcal{L}$ only); if a candidate $c\in\mathcal{L}$ has no mutual neighbor in $\mathcal{L}$, add it to $\mathrm{SC}_{2}$ only if (12) holds.
16:  For each $c\in\mathrm{SC}_{2}$, repeat Lines 7–10 to form groups and append them to $\mathbf{G}$.
17:end if
18:Check overlap: apply the two-stage heuristic in the text (core-based hard assignment, then merge test (13) and (14)).
19:If any centers in $\mathbf{C}$ remain unassigned, assign them by (15).
[…]
In the above procedure, we target a candidate set whose cumulative neighborhood coverage is at least $0.75$. When $\{c_{1},\ldots,c_{T}\}$ yields $\mathcal{P}_{T}<0.75$, the retained local centers may be overly concentrated in overlapping regions, resulting in insufficient coverage. To avoid discarding informative but low-frequency candidates, we expand the candidate set by iteratively adding dropped centers from the initial filtering step (i.e., those with $f=1$ or $r_{i}(\mathbf{x}_{i})>2$) until the target coverage is reached. Details are summarized in Algorithm 2.
Algorithm 2 Local-center filtering for the min strategy
1:Frequency-sorted candidates $\{c_{1},\ldots,c_{T}\}$, full local-center set $\mathcal{C}_{\mathrm{all}}$, neighborhoods $\{\mathbf{N}_{c}\}$ (size $s$), AND scores $\mathrm{AND}(\cdot)$, and sample size $n$.
2:Grouping results $\mathbf{G}$ and retained centers $\mathbf{C}$.
3:Compute $\mathcal{P}_{\mathrm{all}}=\mathcal{P}(\mathcal{C}_{\mathrm{all}})$ and set $\tau\leftarrow\min\{0.75,\;0.9\,\mathcal{P}_{\mathrm{all}}\}$.
4:if $\mathcal{P}(\mathbf{C})<\tau$ then
5:  Initialize $\mathbf{C}\leftarrow\{c_{1},\ldots,c_{T}\}$. and $\mathcal{N}\leftarrow\bigcup_{c\in\mathbf{C}}\mathbf{N}_{c}$.
6:  while $\mathcal{P}(\mathbf{C})<\tau$ do
7:    For each $u\in\mathcal{C}_{\mathrm{all}}\setminus\mathbf{C}$, compute $\Delta(u)=|\mathbf{N}_{u}\setminus\mathcal{N}|$.
8:    Select $u^{*}=\arg\max_{u}\ \Delta(u)\cdot\mathrm{AND}(u)$.
9:    Update $\mathbf{C}\leftarrow\mathbf{C}\cup\{u^{*}\}$ and $\mathcal{N}\leftarrow\mathcal{N}\cup\mathbf{N}_{u^{*}}$.
10:  end while
11:  Apply Algorithm 1 with the candidate set fixed to $\mathbf{C}$ (i.e., skip frequency cutoff stage).
12:else
13:  Apply Algorithm 1 on $\{c_{1},\ldots,c_{T}\}$.
@@
After Algorithm 2, we obtain groups of filtered centers, denoted by $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{\hat{K}}\}$. If any centers in $\mathbf{G}$ violate the min-strategy conditions (7) or (8), we apply the group trimming process in Algorithm 3. The number of clusters is not required; when a user specifies $K$, Algorithm 3 first adjusts the number of groups to $K$ by either dropping less reliable groups (when $\hat{K}>K$) or splitting a weakly cohesive group (when $\hat{K}<K$). For notational simplicity, we denote the left- and right-hand sides of (7) by $\mathrm{mw}(c)$ and $\mathrm{mb}(c)$, respectively.
@@
1:Groups of filtered centers $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{\hat{K}}\}$, neighborhood-similarity matrix $\mathbb{M}$, optional $K$, and optional stability scores $\mathrm{AND}(\cdot)$ for centers.
2:Trimmed groups $\mathbf{G}$ satisfying the min strategy.
3:if $K$ is provided then
4:  if $\hat{K}>K$ then
5:    For each group $\mathbf{g}$, compute $t(\mathbf{g})=\min\{t:c_{t}\in\mathbf{g}\}$.
6:    Drop the (GK)(G-K) groups from $\mathbf{G}$ with the largest $t(\mathbf{g})$.
7:  else if $\hat{K}<K$ then
8:    while $\hat{K}<K$ do
9:     Select the weakest group $\mathbf{g}_{g}$ where
$g=\arg\min_{\ell}\{\min_{c_{u},c_{v}\in\mathbf{g}_{\ell}}\mathbb{M}_{u,v}\}.$
10:     Choose the first prototype center $p\in\mathbf{g}_{g}$ as the most stable one (i.e., it has the highest $\mathrm{AND}$ value within $\mathbf{g}_{g}$).
11:     Choose the second prototype $q\in\mathbf{g}_{g}\setminus\{p\}$ that maximizes
$\left\{\min_{c_{u},c_{v}\in\mathbf{g}^{(1)}(q^{\prime})}\mathbb{M}_{u,v},\ \min_{c_{u},c_{v}\in\mathbf{g}^{(2)}(q^{\prime})}\mathbb{M}_{u,v}\right\},$
where $\mathbf{g}^{(1)}(q^{\prime})$ and $\mathbf{g}^{(2)}(q^{\prime})$ are obtained by assigning each $c\in\mathbf{g}_{g}$ to $\{p,q^{\prime}\}$ with larger neighborhood similarity.
12:     Replace $\mathbf{g}_{g}$ by the two split groups induced by $\{p,q\}$.
13:     Update $\mathbf{G}$, set $\hat{K}\leftarrow\hat{K}+1$.
14:    end while
15:  end if
16:end if
17:while violations of (7) or (8) exist do
18:  while violations of (7) exist do
19:    for each group $\mathbf{g}_{l}\in\mathbf{G}$ with violations of (7) do
20:     Drop the center $c\in\mathbf{g}_{l}$ that maximizes
$\frac{1}{|\mathbf{g}_{l}\setminus\{c\}|}\sum_{c^{\prime}\in\mathbf{g}_{l}\setminus\{c\}}\bigl(\mathrm{mw}(c^{\prime})-\mathrm{mb}(c^{\prime})\bigr).$
21:    end for
22:    Remove any empty groups and update $\mathbf{G}$.
23:  end while
24:  if violations of (8) exist then
25:    Drop all centers violating (8) and update $\mathbf{G}$.
26:  end if
27:end while